Solve the following problems.
In a certain sample, 12% of the original number of radioactive nuclei have undergone disintegration in
a period of 100 years.
1. What percentage of the original radioactive nuclei will remain after 1,000 years?
2. In how many years will only one-fourth of the original number remain?
1
Expert's answer
2021-09-29T04:38:24-0400

(1)

To have less than one left, one needs a DF < 1/1000. no. 12%

This happens after 10 half-lives because (1/2)10 = 1/1024, so after 200 minutes (3 hrs 20 min).


(2)

• Ratio of total radioisotope activity to total mass of the element present

• Measured in Bq/g or Bq/mole

• Useful if radioactive and non-radioactive isotopes of the same element are in mixture (sample with carrier)

• Maximum Specific Activity in carrier-free samples:

Carrier Free Specific Activity (CFSA)


"CFSA [Bq/g] ≈ 4.8\times1018 /(A\times T1/2) for T1/2 \space given\space in\space days,"